Michelle wants to buy a scarf that is on sale for 30% off. She must also pay 6% sales tax. If she buys an item that normally sells for $22.00 plus tax, what is her total cost?
Answer by stanbon(75887)

Michelle wants to buy a scarf that is on sale for 30% off. She must also pay 6% sales tax. If she buys an item that normally sells for $22.00 plus tax, what is her total cost?
----------------
Start $22
Discounted = 0.7*22 = $15.40
Sales tax:: 0.06*15.4 = 92 cents
Ans: Total cost = 15.40+0.92 = $16.32
-------------
Cheers,
Stan H.
-----------------------